# AIM-SLAM: Dense Monocular SLAM via Adaptive and Informative Multi-View Keyframe Prioritization with Foundation Model

### 05 Mar 2026 v1

@@ Abstract
Recent advances in geometric foundation models have emerged as a promising alternative for addressing the challenge of dense reconstruction in monocular visual simultaneous localization and mapping (SLAM). Although geometric foundation models enable SLAM to leverage variable input views, the previous methods remain confined to two-view pairs or fixed-length inputs without sufficient deliberation of geometric context for view selection. To tackle this problem, we propose AIM-SLAM, a dense monocular SLAM framework that exploits an adaptive and informative multi-view keyframe prioritization with dense pointmap predictions from visual geometry grounded transformer (VGGT). Specifically, we introduce the selective information- and geometric-aware multi-view adaptation (SIGMA) module, which employs voxel overlap and information gain to retrieve a candidate set of keyframes and adaptively determine its size. Furthermore, we formulate a joint multi-view $\mathrm{Sim}(3)$ optimization that enforces consistent alignment across selected views, substantially improving pose estimation accuracy. The effectiveness of AIM-SLAM is demonstrated on real-world datasets, where it achieves state-of-the-art performance in both pose estimation and dense reconstruction. Our system supports ROS integration, with code is available at https://aimslam.github.io/.
@@ I Introduction
In this context, we propose AIM-SLAM, a monocular SLAM framework that leverages an adaptive and informative multi-view prioritization with foundation models. By prioritizing informative views for optimization, rather than relying solely on temporally adjacent frames, our approach maximizes geometric consistency, mitigates scale drift, and enables more accurate dense reconstruction.
@@
  • State-of-the-art performance of AIM-SLAM is validated on public datasets, demonstrating improvements in both pose estimation and dense reconstruction. The code is publicly released, with ROS integration also provided.
@@ II-A Classical and Learning-based Visual SLAM
Conventional visual SLAM pipelines could be categorized by their input-processing strategies. Indirect methods estimate motion by minimizing reprojection error of sparse salient primitives, such as points [ptam, vinsmono, orbslam, orbslam3] and line segments [avoiding, plfvins, uvslam]. Direct methods minimize photometric error over pixel intensities, yielding (semi-)dense maps from monocular [lsdslam, dtam, dso], stereo [zhang2021stereo, stereodso], or RGB-D inputs [kerl2013dense, elasticfusion]. Semi-direct approaches combine both paradigms [svo, forster2016svo]. While these approaches have enabled decades of progress, such methods remain constrained by handcrafted modules and reliance on accurate calibration in monocular settings.
With the advent of deep learning [cnn], these handcrafted modules have increasingly been replaced by learned counterparts. This includes data-driven descriptors and matchers for robust feature tracking [mishchuk2017working, superpoint, superglue], end-to-end networks regressing pose (and often depth) from RGB video [sfmlearner, deepvo, undeepvo, ganvo, iyer2018geometric], and hybrid methods injecting learned priors such as depth [codeslam, simsync, d3vo], semantics [cnnslam], or optical flow [d3vo, zhan2020visual] into geometric back-ends. More recently, differentiable dense bundle adjustment (DBA) has been explored in learning-based SLAM: DROID-SLAM [droidslam] coupled ConvGRU updates with DBA, DPVO [dpvo] introduced a path-based formulation, and GO-SLAM [goslam] extended DBA with full bundle adjustment and an implicit surface model.
@@ III-B Problem Definition
AIM-SLAM is a keyframe-based tracking framework, inspired by MASt3R-SLAM [mast3r-slam], and advances beyond fixed-window designs by introducing (i) adaptive, informative and overlap-aware multi-view prioritization regulated by VGGT predictions, and (ii) a joint $\mathrm{Sim}(3)$ optimization that ensures scalable and consistent reconstruction.
@@ IV-A Datasets and Evaluation Metrics
We evaluate AIM-SLAM on the TUM RGB-D dataset [tum] and the EuRoC MAV dataset [euroc]. TUM RGB-D contains room-scale indoor trajectories with cluttered scenes, while EuRoC features aggressive motions and large viewpoint changes, offering a challenging benchmark for robustness.
Evaluation follows two standard metrics for dense visual SLAM: (i) camera pose estimation accuracy, measured by the RMSE of absolute trajectory error (ATE), and (ii) dense reconstruction quality, measured by accuracy, completion, and chamfer distance as similar to previous works [mast3r-slam, vggtslam, vggtlong]. For dense reconstruction evaluations on the EuRoC dataset, we only use the Vicon Room sequences where ground-truth pointclouds were obtained from laser scans.
@@ IV-C2 Dense Reconstruction
On the TUM RGB-D dataset, AIM-SLAM reconstructs fine object details more accurately than baselines, as shown in the bottom row of Fig. 6. The same multi-view strategy also preserves global consistency in large-scale sequences, as demonstrated on the EuRoC dataset (top row of Fig. 6). Here, baseline methods often suffer from ghosting artifacts on planar surfaces due to scale inconsistency, which persist even after two-view $\mathrm{Sim}(3)$ optimization [mast3r-slam] or accumulated submap alignment errors [vggtlong]. These qualitative improvements are corroborated in Table III.
@@
VGGT-Long [vggtlong] 0.137 0.025 0.081 0.102 0.031 0.066
VGGT-SLAM [vggtslam] 0.256 0.033 0.145 0.078 0.030 0.054
MASt3R-SLAM [mast3r-slam] 0.108 0.072 0.090 0.097 0.031 0.064
AIM-SLAM (ours) 0.072 0.031 0.051 0.050 0.026 0.038
@@ V Conclusions
In summary, we presented AIM-SLAM, a dense monocular SLAM framework for uncalibrated settings that leverages geometry-aware foundation models. The proposed SIGMA module adaptively prioritizes a sparse but overlap-rich and informative keyframe subset. The prioritized multi-view subset is then jointly optimized in $\mathrm{Sim}(3)$ space to reduce both short- and mid-term drift. Together, these components enable accurate and globally consistent dense reconstruction under uncalibrated conditions, offering a more scalable solution for foundation model-based SLAM. The current limitation of AIM-SLAM is its reliance on VGGT inference, yielding an overall runtime of about 3 Hz in our environment. Excluding VGGT inference, the remaining components of our method run at about 17 Hz. Future work will explore accelerating the current foundation model or integrating faster alternatives.

### 28 Mar 2026 v3

@@ Abstract
Recent advances in geometric foundation models have emerged as a promising alternative for addressing the challenge of dense reconstruction in monocular visual simultaneous localization and mapping (SLAM). Although geometric foundation models enable SLAM to leverage variable input views, the previous methods remain confined to two-view pairs or fixed-length inputs without sufficient deliberation of geometric context for view selection. To tackle this problem, we propose AIM-SLAM, a dense monocular SLAM framework that exploits an adaptive and informative multi-view keyframe prioritization with dense pointmap predictions from visual geometry grounded transformer (VGGT). Specifically, we introduce the selective information- and geometric-aware multi-view adaptation (SIGMA) module, which employs voxel overlap and information gain to retrieve a candidate set of keyframes and adaptively determine its size. Furthermore, we formulate a joint multi-view $\mathrm{Sim}(3)$ optimization that enforces consistent alignment across selected views, substantially improving pose estimation accuracy. The effectiveness of AIM-SLAM is demonstrated on real-world datasets, where it achieves state-of-the-art pose estimation performance and accurate dense reconstruction results. Our system supports ROS integration, with code is available at https://aimslam.github.io/.
@@ I Introduction
In this context, we propose AIM-SLAM, a monocular SLAM framework that leverages an adaptive and informative multi-view prioritization with foundation models. By prioritizing informative views for optimization, rather than relying solely on temporally adjacent frames, our approach improves geometric consistency, mitigates scale drift, and enables consistent dense reconstruction across diverse scenes.
@@
  • Extensive evaluations on public datasets validate the effectiveness of AIM-SLAM in pose estimation and dense reconstruction. The code is publicly released, with ROS integration also provided.
@@ II-A Classical and Learning-based Visual SLAM
Conventional visual SLAM pipelines could be categorized by their input-processing strategies. Indirect methods estimate motion by minimizing reprojection error of sparse salient primitives, such as points [vinsmono, orbslam, orbslam3] and line segments [avoiding, plfvins, uvslam]. Direct methods minimize photometric error over pixel intensities, yielding (semi-)dense maps from monocular [lsdslam, dtam, dso], stereo [zhang2021stereo, stereodso], or RGB-D inputs [kerl2013dense, elasticfusion]. Semi-direct approaches combine both paradigms [svo, forster2016svo]. While these approaches have enabled decades of progress, such methods remain constrained by handcrafted modules and reliance on accurate calibration in monocular settings.
With the advent of deep learning [cnn], these handcrafted modules have increasingly been replaced by learned counterparts. This includes data-driven descriptors and matchers for robust feature tracking [superpoint, superglue], end-to-end networks regressing pose (and often depth) from RGB video [sfmlearner, deepvo, undeepvo, iyer2018geometric], and hybrid methods injecting learned priors such as depth [codeslam, d3vo], semantics [cnnslam], or optical flow [d3vo, zhan2020visual] into geometric back-ends. More recently, differentiable dense bundle adjustment (DBA) has been explored in learning-based SLAM: DROID-SLAM [droidslam] coupled ConvGRU updates with DBA, DPVO [dpvo] introduced a path-based formulation, and GO-SLAM [goslam] extended DBA with full bundle adjustment and an implicit surface model.
@@ III-B Problem Definition
AIM-SLAM is a keyframe-based tracking framework following MASt3R-SLAM [mast3r-slam], but advances beyond fixed-window designs by introducing (i) adaptive, informative and overlap-aware multi-view prioritization regulated by VGGT predictions, and (ii) a joint $\mathrm{Sim}(3)$ optimization that ensures scalable and consistent reconstruction.
@@ IV-A Datasets and Evaluation Metrics
We evaluate AIM-SLAM on the TUM RGB-D dataset [tum] and the EuRoC MAV dataset [euroc]. TUM RGB-D contains room-scale indoor trajectories with cluttered scenes. EuRoC features aggressive motions and large viewpoint changes, offering a challenging benchmark for robustness; in the uncalibrated setting on EuRoC, undistorted images are used for all methods without providing calibration, following [mast3r-slam].
Evaluation follows two standard metrics for dense visual SLAM: (i) camera pose estimation accuracy, measured by the RMSE of absolute trajectory error (ATE), and (ii) dense reconstruction quality, measured by accuracy, completion, and chamfer distance as similar to previous works [mast3r-slam, vggtslam, vggtlong]. For dense reconstruction evaluations on the EuRoC dataset, we only use the Vicon Room sequences where ground truth pointclouds were obtained from laser scans.
@@ IV-C2 Dense Reconstruction
On the TUM RGB-D dataset, AIM-SLAM reconstructs fine object details more accurately than baselines, as shown in the bottom row of Fig. 6, achieving the strongest reconstruction performance across all metrics. The same multi-view strategy also preserves global consistency in large-scale sequences, as demonstrated on the EuRoC dataset (top row of Fig. 6). Here, baseline methods often suffer from ghosting artifacts on planar surfaces due to scale inconsistency, which persist even after two-view $\mathrm{Sim}(3)$ optimization [mast3r-slam] or accumulated submap alignment errors [vggtlong]. Consistent with this observation, Table III shows that AIM-SLAM achieves the best accuracy and competitive completion and chamfer distance. We attribute this behavior to a trade-off between stable informative-view selection and dense surface coverage, as AIM-SLAM focuses on a compact subset of reliable and overlapping views, which improves pose stability and geometric accuracy.
@@
VGGT-Long [vggtlong] 0.106 0.119 0.112 0.094 0.107 0.100
VGGT-SLAM [vggtslam] 0.246 0.216 0.231 0.109 0.127 0.118
MASt3R-SLAM [mast3r-slam] 0.108 0.072 0.090 0.097 0.113 0.105
AIM-SLAM (ours) 0.103 0.102 0.102 0.063 0.098 0.081
@@ V Conclusions
In summary, we presented AIM-SLAM, a dense monocular SLAM framework for uncalibrated settings that leverages geometry-aware foundation models. The proposed SIGMA module adaptively prioritizes a sparse but overlap-rich and informative keyframe subset. The prioritized multi-view subset is then jointly optimized in $\mathrm{Sim}(3)$ space to reduce both short- and mid-term drift. Together, these components enable accurate pose estimation and geometrically consistent dense reconstruction under uncalibrated conditions, offering a more scalable solution for foundation model-based SLAM. The current limitation of AIM-SLAM is its reliance on VGGT inference, yielding an overall runtime of about 3 Hz in our environment. Excluding VGGT inference, the remaining components of our method run at about 17 Hz. Future work will explore accelerating the current foundation model or integrating faster alternatives.